# Navigating Spatio-Temporal Heterogeneity: A Graph Transformer Approach for Traffic Forecasting

### v1

@@ Abstract.
Traffic forecasting has emerged as a crucial research area in the development of smart cities. Although various neural networks with intricate architectures have been developed to address this problem, they still face two key challenges: i) Recent advancements in network designs for modeling spatio-temporal correlations are starting to see diminishing returns in performance enhancements. ii) Additionally, most models do not account for the spatio-temporal heterogeneity inherent in traffic data, i.e., traffic distribution varies significantly across different regions and traffic flow patterns fluctuate across various time slots. To tackle these challenges, we introduce the Spatio-Temporal Graph Transformer (STGormer), which effectively integrates attribute and structure information inherent in traffic data for learning spatio-temporal correlations, and a mixture-of-experts module for capturing heterogeneity along spaital and temporal axes. Specifically, we design two straightforward yet effective spatial encoding methods based on the graph structure and integrate time position encoding into the vanilla transformer to capture spatio-temporal traffic patterns. Additionally, a mixture-of-experts enhanced feedforward neural network (FNN) module adaptively assigns suitable expert layers to distinct patterns via a spatio-temporal gating network, further improving overall prediction accuracy. Experiments on five real-world datasets demonstrate that STGormer achieves state-of-the-art performance. The code is available at https://github.com/jasonz5/STGormer.
@@ 1. Introduction
To effectively model spatio-temporal correlations and heterogeneity, we propose Spatio-Temporal Graph Transformer (STGormer) for traffic forecasting, and achieves state-of-the-art performance on five real-world traffic datasets. To model spatial and temporal correlations without relying on complex architectures, our STGormer is built over a Transformer model which incorporates several attribute and structure embeddings. Specifically, we add temporal positional information and degree centrality into the node features of the input layer. Additionally, we integrate structural information of the graph into the spatial attention layer, enabling the model to effectively capture spatial and temporal dependencies in traffic data. To address spatial and temporal heterogeneity, our approach includes a mixture-of-experts enhanced FNN module. This module adaptively assigns appropriate expert layers to distinct patterns through a spatio-temporal gating network, which also reveals complex spatio-temporal relationships within the data.
@@
  • Extensive empirical studies conducted on five real-world datasets illustrate that STGormer substantially surpasses baseline models in traffic forecasting accuracy, while also demonstrating superior ability to modeling traffic patterns.
@@ 4. Experiment
METR-LA Traffic Speed 20120301-20120627 Five minutes 207 1515
PEMS-BAY Traffic Speed 20170101–20170630 Five minutes 325 2369
@@ 4.1.1. Datasets and Metrics
To thoroughly evaluate the proposed STGormer model, we conduct extensive experiments on five real-world traffic datasets, as listed in Table 1: NYCTaxi, NYCBike1 (Yao et al., 2019), NYCBike2 (Lin et al., 2019), METR-LA, and PEMS-BAY (Li et al., 2017). These datasets fall into two categories: grid-based and graph-based datasets. For the grid-based datasets (NYCTaxi and NYCBike1/2), we predict the flows for the next time step using the previous 2-hour flows and the previous 3-day flows around the predicted time, ensuring fair comparison with previous work (Ji et al., 2023). For the graph-based datasets (METR-LA and PEMS-BAY), we predict the future 12 frames based on the historical 12 steps. All datasets are divided into training, validation, and test sets with a ratio of 7:1:2.
@@
We implement the model using the PyTorch toolkit on a Linux server equipped with an NVIDIA GeForce RTX 3090 GPU. The embedding dimension $D$ is set to 64. The number of both spatial and temporal MoE-enhanced transformer blocks is set to 1 for the PEMS-BAY dataset and 3 for the other datasets. The number of attention heads is set to 4. Each spatial and temporal router network contains 6 experts. Optimization is performed with Adam optimizer with the learning rate decaying from 0.001, and the batch size is set to 32. We apply an early-stop mechanism if the validation error converges within 25 continuous steps. Hyperparameters were chosen based on commonly used values and adjusted through preliminary experiments, without extensive ablation or hyperparameter optimization.
@@
In this study, we compare our proposed method against several widely used baselines in the field. Historical Average (HA) and Support Vector Regression (SVR) (Smola and Schölkopf, 2004) are traditional statistical methods. For deep learning methods, we include ST-ResNet (Zhang et al., 2017), STGCN (Yu et al., 2017), GMAN (Zheng et al., 2020), Graph WaveNet (Wu et al., 2019), and AGCRN (Bai et al., 2020), each demonstrating significant capabilities in capturing spatial and temporal dependencies within datasets. Additionally, we select STID (Shao et al., 2022b), STEP (Shao et al., 2022a), STAEformer (Liu et al., 2023), and ST-SSL (Ji et al., 2023) as recent state-of-the-art developments to explore advanced neural network architectures and input embedding methods.
@@ 4.2. Main Results
Table 2. Model comparison on five datasets in terms of MAE, RMSE and MAPE (%).
Datasets Metric HA SVR ST-ResNet STGCN GMAN GWNet AGCRN STID STEP STAEformer ST-SSL STGormer
NYCBike1 Horizon 1 (60 min) MAE 7.89 7.63 5.64 5.46 6.97 5.45 5.32 6.66 - 5.14 5.17 4.99
RMSE 12.13 11.35 8.35 8.03 10.11 8.04 7.86 10.33 - 7.51 7.75 7.36
MAPE 28.42 26.41 25.91 27.31 33.23 23.60 26.11 28.26 - 22.86 23.92 23.24
NYCBike2 Horizon 1 (30 min) MAE 12.76 12.15 5.85 5.07 5.11 5.13 4.99 5.39 - 4.84 4.86 4.79
RMSE 18.95 18.59 8.84 7.72 7.84 7.97 7.70 8.58 - 7.24 7.41 7.48
MAPE 45.62 44.22 31.33 26.28 27.07 23.50 26.66 23.89 - 21.72 22.04 21.18
NYCTaxi Horizon 1 (30 min) MAE 48.62 46.94 12.08 11.74 13.58 10.81 11.00 15.56 - 10.72 10.73 10.42
RMSE 83.55 82.98 21.76 21.01 24.85 19.45 20.06 29.97 - 19.13 19.20 19.56
MAPE 65.35 64.58 24.62 20.89 22.35 18.34 18.60 21.10 - 15.94 16.12 15.83
METR-LA Horizon 12 (60 min) MAE 6.99 6.72 - 3.60 3.44 3.51 3.59 3.55 3.37 3.34 3.20 3.10
RMSE 13.89 13.76 - 7.43 7.35 7.28 7.45 7.55 6.99 7.02 6.46 6.34
MAPE 17.54 15.70 - 10.35 10.07 9.96 10.47 10.95 9.61 9.70 8.85 8.39
PEMS-BAY Horizon 12 (60 min) MAE 3.31 3.28 - 2.02 1.92 1.99 1.94 1.91 1.79 1.88 1.81 1.62
RMSE 7.54 7.08 - 4.63 4.49 4.60 4.50 4.42 4.20 4.34 4.13 3.76
MAPE 7.65 8.00 - 4.72 4.52 4.71 4.55 4.55 4.18 4.41 4.31 3.68
  • indicates that the dataset is not suitable for this model or its training configuration.
Table 2 demonstrates that our STGormer framework exhibits enhanced performance across a majority of metrics on all datasets, affirming the efficacy of our approach. For an equitable evaluation, we have sourced the baseline results for the META-LA and PEMS-BAY datasets from the foundational literature, which is highly regarded and repeatedly validated in the realm of spatio-temporal forecasting. Please note that discrepancies in baseline results for the NYCBike1/2 and NYCTaxi datasets stem from our use of ST-SSL’s evaluation methods, specifically calculating prediction errors for values that exceed a non-zero threshold in traffic flow. Other spatio-temporal models also deliver strong results, attributing their success to adept handling of spatio-temporal dependencies. Conversely, traditional non-deep learning techniques such as HA and SVR are the least effective, struggling to model the complex non-linear spatial and temporal dynamics prevalent in real-world datasets. In summary, the STGormer framework significantly advances the state-of-the-art in traffic forecasting, demonstrating its ability to capture the intricate spatio-temporal dynamics.
@@ 4.3. Ablation Study
(a) NYCTaxi
(b) METR-LA
Figure 3. Ablation study.
@@
Figure 3 illustrates the impact of various components on the performance of our model when tested on the NYCTaxi and MATR-LA dataset. The temporal input encoding $t_{in}$ is essential for capturing both periodic and linear aspects of temporal features. Furthermore, the significant performance degradation observed upon removing the spatial input encoding $s_{in}$ and the spatial attention bias matrix $SA_{bias}$ demonstrates that our proposed spatial structure embeddings effectively model the inherent spatial patterns in traffic data and enhance the transformers’ capability to model spatio-temporal graph data. The w/o STMoE variant underscores the crucial role of the Mixture of Experts (MoE) in enhancing the feedforward network’s ability to capture spatio-temporal patterns, where specialized routers direct inputs to the appropriate experts, thereby improving the management of complex traffic patterns. Overall, the results underscore the value of our proposed spatio-temporal graph transformer for traffic forecasting.
@@ 5. Conclusion and Future Work
In this paper, we introduce the Spatio-Temporal Graph Transformer (STGormer) to effectively capture spatio-temporal correlations and heterogeneity in traffic forecasting tasks. Our framework integrates temporal and spatial static context into the analysis of traffic flow data. The resultant fused representation is then fed into spatio-temporal transformers enhanced with Mixture of Experts (MoE), which incorporates spatial context in the attention layer and a spatio-temporal gating network. Finally, leveraging the learned spatio-temporal representation, our framework yields superior prediction results. Extensive experiments conducted on five real-world datasets validate the efficacy of STGormer, outperforming existing traffic forecasting models. In future research, we plan to further enhance and expand the application of STGormer to various spatio-temporal forecasting tasks, extending beyond traffic prediction. Additionally, we aim to explore other spatial structural embedding techniques to improve the transformer’s ability to model spatial relationships, thereby boosting its overall forecasting performance.

### v2

@@ Abstract.
Traffic forecasting has emerged as a crucial research area in the development of smart cities. Although various neural networks with intricate architectures have been developed to address this problem, they still face two key challenges: i) Recent advancements in network designs for modeling spatio-temporal correlations are starting to see diminishing returns in performance enhancements. ii) Additionally, most models do not account for the spatio-temporal heterogeneity inherent in traffic data, i.e., traffic distribution varies significantly across different regions and traffic flow patterns fluctuate across various time slots. To tackle these challenges, we introduce the Spatio-Temporal Graph Transformer (STGormer), which effectively integrates attribute and structure information inherent in traffic data for learning spatio-temporal correlations, and a mixture-of-experts module for capturing heterogeneity along spaital and temporal axes. Specifically, we design two straightforward yet effective spatial encoding methods based on the graph structure and integrate time position encoding into the vanilla transformer to capture spatio-temporal traffic patterns. Additionally, a mixture-of-experts enhanced feedforward neural network (FNN) module adaptively assigns suitable expert layers to distinct patterns via a spatio-temporal gating network, further improving overall prediction accuracy. Experiments on real-world traffic datasets demonstrate that STGormer achieves state-of-the-art performance. The code is available at https://github.com/jasonz5/STGormer.
@@ 1. Introduction
To effectively model spatio-temporal correlations and heterogeneity, we propose Spatio-Temporal Graph Transformer (STGormer) for traffic forecasting, and achieves state-of-the-art performance on three real-world traffic datasets. To model spatial and temporal correlations without relying on complex architectures, our STGormer is built over a Transformer model which incorporates several attribute and structure embeddings. Specifically, we add temporal positional information and degree centrality into the node features of the input layer. Additionally, we integrate structural information of the graph into the spatial attention layer, enabling the model to effectively capture spatial and temporal dependencies in traffic data. To address spatial and temporal heterogeneity, our approach includes a mixture-of-experts enhanced FNN module. This module adaptively assigns appropriate expert layers to distinct patterns through a spatio-temporal gating network, which also reveals complex spatio-temporal relationships within the data.
@@
  • Extensive empirical studies conducted on three real-world datasets illustrate that STGormer substantially surpasses baseline models in traffic forecasting accuracy, while also demonstrating superior ability to modeling traffic patterns.
@@ 4.1.1. Datasets and Metrics
To thoroughly evaluate the proposed STGormer model, we conduct extensive experiments on three real-world traffic datasets, as listed in Table 1: NYCTaxi, NYCBike1 (Yao et al., 2019), NYCBike2 (Lin et al., 2019). For these datasets, we predict the flows for the next time step using the previous 2-hour flows and the previous 3-day flows around the predicted time, ensuring fair comparison with previous work (Ji et al., 2023). All datasets are divided into training, validation, and test sets with a ratio of 7:1:2.
@@
We implement the model using the PyTorch toolkit on a Linux server equipped with an NVIDIA GeForce RTX 3090 GPU. The embedding dimension $D$ is set to 64. The number of both spatial and temporal MoE-enhanced transformer blocks is set to 3. The number of attention heads is set to 4. Each spatial and temporal router network contains 6 experts. Optimization is performed with Adam optimizer with the learning rate decaying from 0.001, and the batch size is set to 32. We apply an early-stop mechanism if the validation error converges within 25 continuous steps. Hyperparameters were chosen based on commonly used values and adjusted through preliminary experiments, without extensive ablation or hyperparameter optimization.
@@
In this study, we compare our proposed method against several widely used baselines in the field. Historical Average (HA) and Support Vector Regression (SVR) (Smola and Schölkopf, 2004) are traditional statistical methods. For deep learning methods, we include ST-ResNet (Zhang et al., 2017), STGCN (Yu et al., 2017), GMAN (Zheng et al., 2020), Graph WaveNet (Wu et al., 2019), and AGCRN (Bai et al., 2020), each demonstrating significant capabilities in capturing spatial and temporal dependencies within datasets. Additionally, we select STID (Shao et al., 2022b), STAEformer (Liu et al., 2023), and ST-SSL (Ji et al., 2023) as recent state-of-the-art developments to explore advanced neural network architectures and input embedding methods.
@@ 4.2. Main Results
Table 2. Model comparison on traffic datasets in terms of MAE, RMSE and MAPE (%).
Datasets Metric HA SVR ST-ResNet STGCN GMAN GWNet AGCRN STID STAEformer ST-SSL STGormer
NYCBike1 Horizon 1 (60 min) MAE 7.89 7.63 5.64 5.46 6.97 5.45 5.32 6.66 5.14 5.17 4.99
RMSE 12.13 11.35 8.35 8.03 10.11 8.04 7.86 10.33 7.51 7.75 7.36
MAPE 28.42 26.41 25.91 27.31 33.23 23.60 26.11 28.26 22.86 23.92 23.24
NYCBike2 Horizon 1 (30 min) MAE 12.76 12.15 5.85 5.07 5.11 5.13 4.99 5.39 4.84 4.86 4.79
RMSE 18.95 18.59 8.84 7.72 7.84 7.97 7.70 8.58 7.24 7.41 7.48
MAPE 45.62 44.22 31.33 26.28 27.07 23.50 26.66 23.89 21.72 22.04 21.18
NYCTaxi Horizon 1 (30 min) MAE 48.62 46.94 12.08 11.74 13.58 10.81 11.00 15.56 10.72 10.73 10.42
RMSE 83.55 82.98 21.76 21.01 24.85 19.45 20.06 29.97 19.13 19.20 19.56
MAPE 65.35 64.58 24.62 20.89 22.35 18.34 18.60 21.10 15.94 16.12 15.83
Table 2 demonstrates that our STGormer framework exhibits enhanced performance across a majority of metrics on all datasets, affirming the efficacy of our approach. Please note that discrepancies in baseline results for the NYCBike1/2 and NYCTaxi datasets stem from our use of ST-SSL’s evaluation methods, specifically calculating prediction errors for values that exceed a non-zero threshold in traffic flow. Other spatio-temporal models also deliver strong results, attributing their success to adept handling of spatio-temporal dependencies. Conversely, traditional non-deep learning techniques such as HA and SVR are the least effective, struggling to model the complex non-linear spatial and temporal dynamics prevalent in real-world datasets. In summary, the STGormer framework significantly advances the state-of-the-art in traffic forecasting, demonstrating its ability to capture the intricate spatio-temporal dynamics.
@@ 4.3. Ablation Study
Figure 3. Ablation study on NYCTaxi.
@@
Figure 3 illustrates the impact of various components on the performance of our model when tested on the NYCTaxi dataset. The temporal input encoding $t_{in}$ is essential for capturing both periodic and linear aspects of temporal features. Furthermore, the significant performance degradation observed upon removing the spatial input encoding $s_{in}$ and the spatial attention bias matrix $SA_{bias}$ demonstrates that our proposed spatial structure embeddings effectively model the inherent spatial patterns in traffic data and enhance the transformers’ capability to model spatio-temporal graph data. The w/o STMoE variant underscores the crucial role of the Mixture of Experts (MoE) in enhancing the feedforward network’s ability to capture spatio-temporal patterns, where specialized routers direct inputs to the appropriate experts, thereby improving the management of complex traffic patterns. Overall, the results underscore the value of our proposed spatio-temporal graph transformer for traffic forecasting.
@@ 5. Conclusion and Future Work
In this paper, we introduce the Spatio-Temporal Graph Transformer (STGormer) to effectively capture spatio-temporal correlations and heterogeneity in traffic forecasting tasks. Our framework integrates temporal and spatial static context into the analysis of traffic flow data. The resultant fused representation is then fed into spatio-temporal transformers enhanced with Mixture of Experts (MoE), which incorporates spatial context in the attention layer and a spatio-temporal gating network. Finally, leveraging the learned spatio-temporal representation, our framework yields superior prediction results. Extensive experiments conducted on real-world traffic datasets validate the efficacy of STGormer, outperforming existing traffic forecasting models. In future research, we plan to further enhance and expand the application of STGormer to various spatio-temporal forecasting tasks, extending beyond traffic prediction. Additionally, we aim to explore other spatial structural embedding techniques to improve the transformer’s ability to model spatial relationships, thereby boosting its overall forecasting performance.